# Vid-LLM: A Compact Video-based 3D Multimodal LLM with Reconstruction–Reasoning Synergy

### v1

@@ 1 INTRODUCTION
To overcome these limitations, a more general solution is to enable the model to directly reconstruct scene geometry from video (Leroy et al., 2024; Wang et al., 2024), thereby eliminating the reliance on external depth, pose, or registration modules. More importantly, reconstruction and reasoning are intrinsically interdependent: geometric structures underpin semantic understanding, while semantic reasoning, in turn, provides contextual priors that guide and refine geometric modeling  (Cheng et al., 2024a; Ha & Song, 2022).
@@ 2 Related Work
3D-aware Multimodal Large Language Models (3D-MLLMs). 3D-MLLMs aim to unify 3D scene understanding and vision–language reasoning within a unified linguistic interface, representing an important extension of multimodal large language models (MLLMs). Existing approaches predominantly rely on explicit geometric inputs: some leverage point clouds or reconstructed scenes, often augmented with rendered views, region-level alignments, or condensed 3D feature grids to support large-scale embodied training (Hong et al., 2023a; Chen et al., 2024b; c; Fu et al., 2024; Huang et al., 2023b); others map 3D features to the language space and model spatial relations to enable interactive dialogue (Huang et al., 2023a; 2024). Despite their progress, these methods invariably depend on complex inputs such as point clouds, reconstructed scenes, multi-view renderings, or object-level annotations, which impose substantial burdens on data acquisition, preprocessing, and computation, thereby limiting scalability and transferability.
3D Reconstruction. 3D reconstruction has evolved from multi-view geometry pipelines to neural implicit representations and, more recently, feed-forward transformer-based architectures. Classical methods yield accurate geometry but require dense views and heavy preprocessing (Schonberger & Frahm, 2016; Furukawa et al., 2015). Neural radiance fields and point-based extensions improve fidelity and efficiency but focus mainly on appearance modeling while lacking semantic reasoning (Mildenhall et al., 2021; Barron et al., 2022; Kerbl et al., 2023). Recent feed-forward approaches enable direct prediction of depth, pose, and point clouds from video inputs  (Wang et al., 2024; Leroy et al., 2024; Wang et al., 2025a). Nevertheless, 3D reconstruction is still only weakly integrated into vision–language research, and its role in supporting semantic reasoning remains underexplored.
@@ 3.4 Training Strategy
where $L^{feat}_{geo}$, $L^{feat}_{lang}$ and $L_{sc}$ are the geometry loss, semantic loss and structural consistency loss, respectively. $\lambda$ is a balancing hyperparameter. The geometry loss and semantic loss are defined as:
$L_{geo}^{feat}=\tfrac{1}{N}\sum_{i=1}^{N}\lVert\mathrm{Norm}(T^{\prime}_{\mathrm{geom},i})-\mathrm{Norm}(T_{tea,i}^{geo})\rVert_{2}^{2},\quad L_{lang}^{feat}=1-\cos(T^{\prime}_{\mathrm{lang}},T_{tea}^{lang})$ (6)
where $T^{\prime}_{geom}$ and $T^{\prime}_{lang}$ are the task‑specific geometric and semantic features, respectively, extracted by the Cross-Task Adapter, serving as the student representations in the distillation strategy. $T_{tea}^{geom}$ and $T_{tea}^{lang}$ are the features of the DINO and CLIP encoder and serve as the teacher representations. $N$ denotes the number of feature tokens sampled for alignment. To maintain structural consistency, we also introduce the structural consistency loss $L_{sc}$, which is defined as:
@@
$Z_{stu}=[\text{Norm}(T_{\text{geom}}^{\prime});\text{Norm}(T_{\text{lang}}^{\prime})]\in\mathbb{R}^{M\times C}$ concatenates the geometry and semantic tokens from student representation. $Z_{tea}=[\text{Norm}(T_{\text{tea}}^{\text{geo}});\text{Norm}(T_{\text{tea}}^{\text{lang}})]$ is similarly defined for the teacher representation. $\|\cdot\|_{F}$ is the Frobenius norm, and $M$ is the total number of tokens.
@@
where $e_{i}=\log(d_{i}^{pred}+\varepsilon)-\log(d_{i}^{gt}+\varepsilon)$ is the log-depth error where $\varepsilon$ is a small constant for numerical stability. $b=\tfrac{1}{K}\sum_{i=1}^{K}e_{i}$ is the mean error across all $K$ valid pixels in the image. The parameter $\alpha>0$ controls the robustness by down-weighting large residuals.
@@ 4 Experiments
This section provides a comprehensive evaluation of Vid-LLM. In Section 4.1, the model is benchmarked on 3D vision–language reasoning tasks against state-of-the-art methods. In Section 4.2, the comparison with naive concatenation pipelines highlighting the importance of feature-level integration over post-hoc combination. In Section 4.3, ablation studies analyze the contributions of core modules. Further implementation details are provided in Appendix A.3.
4.1 3D Vision–Language Reasoning
Overview. To comprehensively assess the performance of Vid-LLM on 3D vision–language reasoning tasks, we conduct experiments using widely adopted datasets covering three task categories: 3D Question Answering on ScanQA (Azuma et al., 2022) and SQA3D (Ma et al., 2022), 3D Dense Captioning on Scan2Cap (Chen et al., 2021), and 3D Visual Grounding on ScanRefer (Chen et al., 2020), Multi3DRefer (Zhang et al., 2023), and Nr3D/Sr3D (Achlioptas et al., 2020b). We follow the standard evaluation protocols and metrics defined for each dataset. For grounding tasks, Vid-LLM directly outputs 3D bounding boxes on ScanRefer and Multi3DRefer, which target scene-level grounding with natural language descriptions. For Nr3D/Sr3D, we adopt the Chat-3D v2 protocol (Huang et al., 2023a). which employs the ViL3DRel (Chen et al., 2022) method to generate grounding candidates, emphasizing instance-level referential resolution.
Baseline. For the 3D Question Answering and 3D Dense Captioning tasks, we compare Vid-LLM against representative 3D vision–language reasoning models. The comparison methods include: ScanQA (Azuma et al., 2022), Scan2Cap (Chen et al., 2021), 3D-VisTA (Zhu et al., 2023), 3D-LLM (Hong et al., 2023b), LL3DA (Chen et al., 2024b), Grounded3D-LLM (Chen et al., 2024c), Chat-3D v2 (Huang et al., 2023a), LEO (Huang et al., 2023b), Scene-LLM (Fu et al., 2024), and ChatScene (Huang et al., 2024). All these methods rely on explicit 3D scene inputs. We also compare recent video-based approaches, namely VILA-40B (Lin et al., 2024), IXC-2.5-7B (Zhang et al., 2024b), LLaVA-OV-7B (Li et al., 2024), LLaVA-Video-7B (LLaVA Team, 2024), and Uni3DR2 (Chu et al., 2024). For the 3D visual grounding benchmarks, we compare Vid-LLM with task-specific grounding models including ScanRefer (Chen et al., 2020), 3D-VisTA (Zhu et al., 2023), ReferIt3D (Achlioptas et al., 2020a), 3DVG-Trans (Zhao et al., 2021), MVT (Huang et al., 2022), ViL3DRel (Chen et al., 2022), and SceneVerse (Jia et al., 2024). We also compare against recent 3D-MLLMs capable of performing grounding tasks, including 3D-LLM (Hong et al., 2023a), Chat-3D v2 (Huang et al., 2023a), Grounded3D-LLM (Chen et al., 2024c), and ChatScene (Huang et al., 2024). These comparisons collectively provide a comprehensive evaluation of our proposed method for 3D vision–language (VL) reasoning.
Table 1: Evaluation of 3D Question Answering on ScanQA and SQA3D. Methods marked with * are 3D MLLM evaluated in video mode. ”C” stands for ”CIDEr”, ”B-4” for ”BLEU-4”, ”M” for ”METEOR”, ”R” for ”ROUGE”, and ”EM@1” for top-1 exact match.
@@
Vid-LLM (Ours) 50.1 46.7 47.2 42.9
Result & Analysis. From Tab. 1, we can observe that Vid-LLM achieves the best performance in almost all metrics on the 3D Question Answering benchmarks, which verifies the effectiveness of our proposed method. Compared with 3D-based models and 2D-based models, our method achieves an average improvement of 9% and 31% over the second-best results, respectively. This illustrates that the integration of geometric cues into sematic reasoning can significantly improve perception performance. Furthermore, we also carried out experiments on the 3D Dense Captioning and 3D visual grounding task, which can be seen in Tab. 2 and Tab. 3. Benefiting from our designed cross-task adapter that compactly integrates semantic and geometric information, our method demonstrates superior multi-task capabilities. It is worth noting that in Tab. 3, our method ranks second to ChatScene. This is because ChatScene employs point cloud inputs with instance segmentation, which require much higher acquisition cost than the video inputs used in our model. For qualitative analysis, we visualize the 3D grounding on the SCanRefer dataset in Fig. 4. The results intuitively illustrates the predictive capability of 3D grounding in typical scenarios.
@@
Table 5: Comparison with Naive Concatenation on Reconstruction, 3D VL, and Efficiency. The reported metrics are averaged over all evaluation scores on ScanNet and ScanQA. Reconstruction is evaluated without scale normalization.
Method Recon 3D VL Tasks Time (s / scene) $\downarrow$
ScanNet $\uparrow$ ScanQA $\uparrow$
VGGT+LLaVA3D 24.6 2.7s
VGGT+real-scale+LLaVA-3D 0.587 40.3 2.7s
VGGT+real-scale+LLaVA-3D* 0.587 39.1 2.7s
VGGT*+real-scale+LLaVA-3D 0.591 42.1 2.7s
VGGT*+real-scale+LLaVA-3D* 0.591 41.6 2.7s
Vid-LLM (Ours) 0.582 45.7 1.6s
4.2 Comparison with Naive Concatenation
Overview. To examine whether video-based 3D vision–language (3D VL) tasks can be addressed by directly combining a reconstruction model with a 3D multimodal LLM, we build a concatenation baseline by feeding the 3D reconstruction results from VGGT (Wang et al., 2025a) into the reasoning model LLaVA-3D (Zhu et al., 2025), since our reconstruction backbone is derived from VGGT and our reasoning branch is based on LLaVA-3D. For a fair comparison, we consider several baseline formulations. First, because the raw outputs of VGGT do not preserve metric scale, we additionally evaluate settings where its predictions are aligned with metric scale before being passed to LLaVA-3D. Second, since our Vid-LLM is fine-tuned on ScanNet and its 3D-LLM module is further fine-tuned on predicted 3D scenes, we correspondingly fine-tune VGGT on ScanNet (denoted VGGT*) and take the point clouds produced by both VGGT and VGGT as geometric supervision to fine-tune LLaVA-3D (denoted LLaVA-3D*).
Result & Analysis. As shown in Tab. 5, compared with other settings that incorporate metric-scale alignment, the direct concatenation of VGGT with LLaVA-3D is largely ineffective for 3D VL tasks, highlighting that the absence of metric-scale cues renders geometric representations inadequate for reasoning. With metric-scale alignment, the 3D VL score increases substantially from 24.6 to 40.3, demonstrating the necessity of metric-scale alignment. A further comparison between VGGT and VGGT* shows that fine-tuning VGGT on ScanNet improves the reconstruction metric from 0.587 to 0.591. This improvement in geometric accuracy also leads to a higher 3D VL score, increasing from 40.3 to 42.1. However, when LLaVA-3D is further fine-tuned using predicted point clouds as geometric supervision, the 3D VL score decreases (42.1 $\rightarrow$ 41.6), suggesting that direct supervision from predicted geometry can introduce geometric bias and weaken semantic consistency. Finally, a comparison of VGGT*+LLaVA-3D* with our Vid-LLM highlights the main advantage of our approach: Vid-LLM achieves a notable improvement in 3D VL reasoning (45.7 vs. 42.1) while also reducing inference time (1.6s vs. 2.7s). It is worth noticing that the reconstruction metric of our method is slightly lower than the best baseline, this is because our model prioritizes geometry-semantics interaction over raw geometric accuracy. Our advantage derives from the compact design and the Cross-Task Adapter, which aligns 3D geometry priors with vision–language representations and provides geometric constraints that improve the robustness of reasoning under geometric uncertainty. By facilitating effective geometry–semantics interaction, CTA ensures that Vid-LLM maintains strong reasoning performance even without relying on precise external 3D inputs. Further analysis of our model’s reconstruction performance can be found in Appendix A.2.
4.3 Ablation Studies
@@
Figure 5: Ablation on Cross-Task adapter (CTA) and Metric Depth (MD) Modules. Results are reported as the mean of evaluation metrics across the corresponding datasets.
@@
Cross-Task adapter. As shown in Tab. 6, we conduct ablation studies on the Cross-Task adapter to examine its role in video-based 3D MLLM. Since 3D VL reasoning relies on 3D scene information provided by geometric reconstruction, both tasks must be jointly optimized to achieve strong performance. Without CTA, neither reconstruction nor reasoning performs well, revealing that a single visual feature representation is insufficient to serve both tasks simultaneously. Introducing CTA brings significant improvements to both reconstruction and reasoning by facilitating cross-task feature interaction. However, without bridge tokens, the adapter cannot effectively align geometry priors with semantic features, resulting in notably weaker performance. Adding bridge tokens leads to significant gains, and using 16 tokens achieves the best trade-off between accuracy and model complexity by ensuring sufficient interaction without redundancy.
Metric Depth Modules. Tab. 6 also reports ablations on the use of metric depth and our alignment strategy. Without metric depth supervision (w/o MD), scale ambiguity destroys geometric consistency and leads to near failure of 3D VL reasoning. Incorporating metric depth without alignment (MD – w/o Alignment) preserves absolute scale but produces less accurate reconstruction, which in turn limits the performance of 3D VL tasks. By contrast, combining metric depth with our scale alignment strategy effectively exploits both global scale cues and relative structural details, yielding the most accurate reconstruction and consistently stronger results on 3D VL tasks.
Two-Stage Training. Fig. 6 compares different training strategies to assess their impact on convergence and performance. The single-stage setting converges slowly and stabilizes at a relatively high loss value, reflecting gradient interference when all modules are trained jointly. The two-stage strategy with a single teacher offers limited supervisory signals, limiting optimization quality. Removing the relational loss also degrades convergence and accuracy, highlighting the importance of cross-task consistency. In contrast, our two-stage strategy converges fastest and achieves the lowest final loss.

### v2

@@ 1 INTRODUCTION
To overcome these limitations, a more general solution is to enable the model to directly reconstruct scene geometry from video (Leroy et al., 2024; Wang et al., 2024), thereby eliminating the reliance on external depth, pose, or registration modules. More importantly, reconstruction and reasoning are intrinsically interdependent: geometric structures underpin semantic understanding, while semantic reasoning, in turn, provides contextual priors that guide and refine geometric modeling  (Cheng et al., 2024a; Ha and Song, 2022).
@@ 2 Related Work
3D-aware Multimodal Large Language Models (3D-MLLMs). 3D-MLLMs aim to unify 3D scene understanding and vision–language reasoning within a unified linguistic interface, representing an important extension of multimodal large language models (MLLMs). Existing approaches predominantly rely on explicit geometric inputs: some leverage point clouds or reconstructed scenes, often augmented with rendered views, region-level alignments, or condensed 3D feature grids to support large-scale embodied training (Hong et al., 2023a; Chen et al., 2024b; c; Fu et al., 2024; Huang et al., 2023b); others map 3D features to the language space and model spatial relations to enable interactive dialogue (Huang et al., 2023a; 2024; Zheng et al., 2025b; Huang et al., 2025). Despite their progress, these methods invariably depend on complex inputs such as point clouds, reconstructed scenes, multi-view renderings, or object-level annotations, which impose substantial burdens on data acquisition, preprocessing, and computation, thereby limiting scalability and transferability. In addition, the recent approach VGLLM (Zheng et al., 2025a) explores a video-based 3D-MLLM setting by adopting the geometry encoder of VGGT (Wang et al., 2025a) to extract 3D geometry features from video.
3D Reconstruction. 3D reconstruction has evolved from multi-view geometry pipelines to neural implicit representations and, more recently, feed-forward transformer-based architectures. Classical methods yield accurate geometry but require dense views and heavy preprocessing (Schonberger and Frahm, 2016; Furukawa et al., 2015). Neural radiance fields and point-based extensions improve fidelity and efficiency but focus mainly on appearance modeling while lacking semantic reasoning (Mildenhall et al., 2021; Barron et al., 2022; Kerbl et al., 2023). Recent feed-forward approaches enable direct prediction of depth, pose, and point clouds from video inputs  (Wang et al., 2024; Leroy et al., 2024; Wang et al., 2025a). Nevertheless, 3D reconstruction is still only weakly integrated into vision–language research, and its role in supporting semantic reasoning remains underexplored.
@@ 3.4 Training Strategy
where $L^{feat}_{geo}$, $L^{feat}_{lang}$ and $L_{sc}$ are the geometry loss, semantic loss and structural consistency loss, respectively. $\lambda$ is a balancing hyperparameter. For geometry learning, an L2 loss is applied to patch-level features, following DINO’s patch regression strategy. For semantic learning, a cosine similarity loss is applied to pooled features to align with CLIP’s global semantic embedding space. The specific forms are defined as follows:
$L_{geo}^{feat}=\tfrac{1}{N}\sum_{i=1}^{N}\lVert\mathrm{Norm}(T^{\prime}_{\mathrm{geom},i})-\mathrm{Norm}(T_{tea,i}^{geo})\rVert_{2}^{2},\quad L_{lang}^{feat}=1-\cos(\mathrm{Pool}(T^{\prime}_{\mathrm{lang}}),\mathrm{Pool}(T_{tea}^{lang}))$ (6)
where $\mathrm{Norm}(\cdot)$ denotes L2 normalization, and $\mathrm{Pool}(\cdot)$ indicates mean pooling. $T^{\prime}_{geom}$ and $T^{\prime}_{lang}$ are the task-specific geometric and semantic features extracted by the Cross-Task Adapter, and serve as the student representations in the distillation strategy. $T_{tea}^{geom}$ and $T_{tea}^{lang}$ are features from the DINO and CLIP encoders and serve as teacher representations. $N$ represents the number of tokens in $T^{\prime}_{geom}$. Before computing $L^{feat}_{geo}$ and $L^{feat}_{lang}$, we apply lightweight linear projection layers to align the features to the same embedding dimension. To maintain structural consistency, we also introduce the structural consistency loss $L_{sc}$, which is defined as:
@@
$Z_{stu}=[\text{Norm}(T_{\text{geom}}^{\prime});\text{Norm}(T_{\text{lang}}^{\prime})]\in\mathbb{R}^{M\times C}$ is obtained by concatenating the geometry and semantic tokens from the student representation. $Z_{tea}=[\text{Norm}(T_{\text{tea}}^{\text{geo}});\text{Norm}(T_{\text{tea}}^{\text{lang}})]$ is defined in the same way as $Z_{stu}$, but using the teacher representation. $\|\cdot\|_{F}$ is the Frobenius norm, $M$ is the total number of tokens, and $C$ is the embedding dimension of each token.
@@
The log-depth error is defined as $e_{i}=\log(d_{i}^{pred}+\varepsilon)-\log(d_{i}^{gt}+\varepsilon)$, where $\varepsilon$ is a small constant for numerical stability. $b=\tfrac{1}{K}\sum_{i=1}^{K}e_{i}$ is the mean error across all $K$ valid pixels in the image. The parameter $\alpha>0$ controls the robustness by down-weighting large residuals.
During joint training, $L_{recon-task}$ and $L_{MD}$ optimize the 3D reconstruction model, $L_{VL-task}$ optimize the 3D vision-language model, while the shared CTA is jointly optimized by $L_{recon-task}$ and $L_{VL-task}$. It is worth to noticing that the loss constructed in the 3D-VL reasoning branch is not used to optimize the 3D reconstruction model, which can be seen in Fig. 2. This one-way gradient flow ensures that the CTA acts as a stable bridge for geometry-semantics interaction, enabling effective feature exchange.
@@ 4 Experiments
This section provides a comprehensive evaluation of Vid-LLM. In Section 4.1, we introduce the experimental setup, covering the training details of Vid-LLM and the evaluation settings used in our experiments. In Section 4.2, we evaluate the model on 3D vision–language reasoning tasks against state-of-the-art methods. In Section 4.3, we compare Vid-LLM with representative joint reconstruction–reasoning models to examine how different design choices affect performance when jointly executing reconstruction and 3D VL reasoning. In Section 4.4, we conduct ablation studies to analyze the contributions of core modules.
4.1 Experimental Setup
Training Details. We adopt DINOv2-L as the visual backbone, consisting of 24 Transformer layers with a hidden dimension of 1024. The projection MLPs use two fully connected layers with an expansion factor of 4 and GELU activation. For data processing, we uniformly sample 32 frames per scene, resize the shorter side to 518, crop the resolution so that both height and width are multiples of 14. Optimization is performed using AdamW with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and weight decay of $0.05$, with gradient clipping at $1.0$.
Evaluation Details. We follow the standard evaluation protocols and metrics defined for each dataset. For tables that summarize performance across multiple datasets (e.g., Tables 4.3 and 6), the reported numbers are computed as the mean of the metrics for each benchmark. For the Scan2Cap dataset, we follow prior work (Zhu et al., 2025; Zheng et al., 2025a): instance proposals are generated using Mask3D (Schult et al., 2022), and 3D coordinate tokens are constructed from the predicted instance centers to enable instance-aware caption generation. For the ScanRefer dataset, Mask3D (Schult et al., 2022) serves as the 3D segmentor to remain consistent with supervised baselines (Huang et al., 2024; Chen et al., 2024c; Huang et al., 2023a). For the Nr3D and Sr3D datasets, the provided segmentation annotations are adopted to align with previous methods (Chen et al., 2022; Huang et al., 2023a). For reconstruction evaluation on ScanNet, the metrics are computed in metric scale. All models are trained with the same data and settings for a fair comparison.
Table 1: Evaluation of 3D Question Answering on ScanQA and SQA3D. Methods marked with * are 3D MLLM evaluated in video mode. indicates the model consumes VGGT-generated 3D geometry. ”C” stands for ”CIDEr”, ”B-4” for ”BLEU-4”, ”M” for ”METEOR”, ”R” for ”ROUGE”, and ”EM@1” for top-1 exact match.
@@ 4.1 Experimental Setup
Video-3D LLM 58.1 51.7 58.0 52.7
3DRS 62.9 56.1 60.4 54.9
[…]
VGLLM 53.5 47.5
3DRS 55.2 51.1 52.8 48.3
Vid-LLM (Ours) 63.2 56.4 61.6 56.1
4.2 3D Vision–Language Reasoning
Overview. To comprehensively assess the performance of Vid-LLM on 3D vision–language reasoning tasks, we conduct experiments using widely adopted datasets covering three task categories: 3D Question Answering on ScanQA (Azuma et al., 2022) and SQA3D (Ma et al., 2022), 3D Dense Captioning on Scan2Cap (Chen et al., 2021), and 3D Visual Grounding on ScanRefer (Chen et al., 2020), Multi3DRefer (Zhang et al., 2023), and Nr3D/Sr3D (Achlioptas et al., 2020b).
Baseline. We compare Vid-LLM with a broad set of 3D-based and video-based vision–language reasoning models. 3D-based methods rely on explicit 3D scene inputs such as point clouds or reconstructed geometry, whereas video-based methods operate solely on video. The 3D-based baselines include ScanQA (Azuma et al., 2022), Scan2Cap (Chen et al., 2021), 3D-VisTA (Zhu et al., 2023), 3D-LLM (Hong et al., 2023b), LL3DA (Chen et al., 2024b), Grounded3D-LLM (Chen et al., 2024c), Chat-3D v2 (Huang et al., 2023a), LEO (Huang et al., 2023b), Scene-LLM (Fu et al., 2024), ChatScene (Huang et al., 2024), LLaVA-3D (Zhu et al., 2025), Video-3D LLM (Zheng et al., 2025b), and 3DRS (Huang et al., 2025). The video-based baselines include VILA-40B (Lin et al., 2024), IXC (Zhang et al., 2024b), LLaVA-OV (Li and others, 2024), LLaVA-Video (LLaVA Team, 2024), Uni3DR2 (Chu et al., 2024), and VGLLM (Zheng et al., 2025a). To further assess the grounding capability of Vid-LLM, we additionally compare against task-specific 3D visual grounding models, including ScanRefer (Chen et al., 2020), 3D-VisTA (Zhu et al., 2023), ReferIt3D (Achlioptas et al., 2020a), 3DVG-Trans (Zhao et al., 2021), MVT (Huang et al., 2022), ViL3DRel (Chen et al., 2022), and SceneVerse (Jia et al., 2024). To allow comparison under the same video input setting with the most competitive 3D-based baseline in our evaluation, we additionally include 3DRS, a variant of 3DRS that takes VGGT-generated 3D geometry as input. These comparisons collectively provide a comprehensive evaluation of Vid-LLM across 3D vision–language reasoning tasks.
Result & Analysis. Vid-LLM consistently shows robust performance across all 3D vision–language reasoning benchmarks. As shown in Tab. 1, Vid-LLM achieves the best results among video-based models on both ScanQA and SQA3D, outperforming the second-best baseline (3DRS) by an average margin of 11%. Tab. 2 further shows that Vid-LLM attains the highest performance among video-based methods on Scan2Cap; notably, its M@0.5 score (28.7) is close to that of the best-performing 3D-based model (29.0), despite not using depth or point clouds. These results collectively indicate that the integration of geometric cues into semantic reasoning can significantly improve perception performance. Tab. 3 and Tab. 4.3 present results on ScanRefer/Multi3DRefer and Nr3D/Sr3D. Vid-LLM achieves the best performance across all metrics on the two grounding benchmarks, surpassing all comparable 3D-based and video-based counterparts. These results demonstrate that the proposed Cross-Task Adapter, which effectively integrates semantic and geometric information, enables effective spatial reasoning and yields substantial gains on geometry-intensive tasks such as 3D visual grounding. For qualitative analysis, we visualize the 3D grounding results on the ScanRefer dataset in Fig. 4, which further demonstrates the 3D visual grounding capability of Vid-LLM. More qualitative visualizations for 3D VL tasks can be found in Appendix A.1.
4.3 Comparison with Joint Reconstruction–Reasoning Models
Overview. In addition to 3D vision–language reasoning, Vid-LLM also incorporates a reconstruction branch that provides geometric priors and metric-scale cues. This naturally raises the question of how different design choices for coupling reconstruction and reasoning affect performance when both tasks must be performed from a single video input. To provide a comprehensive analysis, we compare Vid-LLM with two representative categories of joint reconstruction–reasoning baselines: (i) simple concatenation pipelines that feed the 3D geometry reconstructed from video into a 3D multimodal LLM, and (ii) end-to-end architectures that integrate reconstruction and 3D VL reasoning within a single model.
@@ 4.3 Comparison with Joint Reconstruction–Reasoning Models
Table 5: Comparison with end-to-end frameworks for joint 3D reconstruction and 3D VL reasoning.
[…]
VGGT+LLaVA-3D 24.6 35.1 31.6 2.7
VGGT+LLaVA-3D 0.591 42.1 50.9 47.3 2.7
VGGT+LLaVA-3D†‡ 0.591 41.6 48.2 45.9 2.7
[…]
Uni3DR2 0.580 30.2 42.7 33.6 2.1
VGLLM-Rec 0.541 40.9 51.6 48.5 1.8
Vid-LLM (Ours) 0.582 45.7 53.2 59.8 1.6
  • $\dagger$
    $\dagger$ denotes introducing real-scale;
  • $\ddagger$
    $\ddagger$ denotes training LLaVA-3D with 3D geometry predicted by VGGT
Baseline. For the concatenation baselines, we construct pipelines that feed the predicted geometry information of VGGT into LLaVA-3D (Zhu et al., 2025) for 3D vision–language reasoning. Three variants of the concatenation pipeline are evaluated: (i) VGGT+LLaVA-3D, the direct combination of VGGT and LLaVA-3D; (ii) VGGT+LLaVA-3D, which introduces metric-scale alignment to the predictions of VGGT; (iii) VGGT+LLaVA-3D†‡, which further fine-tunes LLaVA-3D using depth and camera poses predicted by VGGT as geometric supervision. For end-to-end joint-task baselines, we include Uni3DR2 (Chu et al., 2024), an end-to-end framework that integrates a reconstruction module with a 3D VL reasoning branch. Notably, Uni3DR2 requires ground-truth camera poses as input, whereas Vid-LLM performs both tasks directly from video, making the comparison conservative in favor of Uni3DR2. We also include VGLLM-Rec, an extension of VGLLM (Zheng et al., 2025a). To support joint reconstruction and reasoning, we reconnect the original camera and depth prediction heads of VGGT to the geometry encoder in VGLLM, enabling VGLLM-Rec to generate geometric predictions in addition to performing 3D VL reasoning from video.
Result & Analysis. As shown in Tab. 4.3, VGGT+LLaVA-3D obtains very low 3D VL accuracy, indicating that relative-scale geometry cannot provide reliable cues for 3D VL reasoning. Building on this, VGGT+LLaVA-3D shows a clear performance gain once metric-scale alignment is applied. However, even with this setting, the concatenation baseline still underperforms Vid-LLM on all 3D VL tasks despite achieving a higher reconstruction score (0.591 vs. 0.582). This indicates that simply providing accurate reconstructed geometry to a 3D-LLM is insufficient and that effective geometry–semantics interaction is necessary for reliable reasoning performance. Furthermore, VGGT+LLaVA-3D†‡ yields a slight drop in 3D VL accuracy compared with VGGT+LLaVA-3D, suggesting that training the 3D-LLM with noisy predicted geometry may introduce biases that degrade the semantic representations of the framework. In contrast to concatenation pipelines, end-to-end joint-task models avoid multi-stage processing and achieve lower inference latency, reducing runtime from 2.7 s/scene to 1.6–2.1 s/scene. From Tab. 4.3, we can observe that Vid-LLM achieves the best overall performance among the joint models on both reconstruction and 3D VL tasks. It marginally outperforms Uni3DR² in reconstruction quality on ScanNet (0.582 vs. 0.580) despite Uni3DR² having access to ground-truth camera poses. On the 3D VL tasks, Vid-LLM demonstrates a clear advantage, with an average relative improvement of 47.6% over Uni3DR² and 12.1% over VGLLM-Rec across ScanQA, Scan2Cap, and ScanRefer. Overall, Vid-LLM delivers the fastest inference speed and the best 3D VL performance, while also achieving the best reconstruction quality among joint-task models. These gains stem from our integrated architecture and the Cross-Task Adapter, which facilitates geometry–semantics interaction and allows the model to effectively leverage geometric cues during joint reasoning. More experimental results about reconstruction performance can be found in Appendix A.2.
@@
Figure 5: Ablation on Cross-Task Adapter (CTA) and Metric Depth (MD) Modules.
@@
4.4 Ablation Studies
Cross-Task Adapter. To assess the contribution of the Cross-Task Adapter, we compare several configurations: removing the CTA module (w/o CTA); in the setting without bridge tokens, applying self-attention to the concatenated $T_{\text{geom}}$ and $T_{\text{lang}}$ (CTA–Concat-SA) and performing cross-attention between the two feature sets (CTA–w/o Bridge (CA)); and finally using the full CTA with bridge tokens. We also analyze how different numbers of bridge tokens (4, 8, 16, and 32) affect model performance. As shown in Tab. 6, w/o CTA leads to the lowest performance on both tasks, confirming that a shared visual representation alone cannot jointly support reconstruction and 3D VL reasoning. Both CTA–Concat-SA and CTA–w/o Bridge(CA) show noticeable improvements, yet their effectiveness remains limited because neither design provides a stable shared latent space for passing geometric cues to semantic features. The configurations using the full CTA module (CTA–4/8/16/32tokens) all achieve clear performance improvements, since the bridge tokens provide a dedicated latent space that enables consistent geometry–semantic alignment and more effective cross-task interaction. Among different token counts, using 16 bridge tokens achieves the best balance between accuracy and model complexity.
Metric Depth Modules. Tab. 6 also reports ablations on the use of metric depth and our alignment strategy. Without metric depth supervision (w/o MD), scale ambiguity destroys geometric consistency and leads to near failure of 3D VL reasoning. Incorporating metric depth without alignment (MD – w/o Alignment) preserves metric scale but produces less accurate reconstruction, which in turn limits the performance of 3D VL tasks. By contrast, combining metric depth with our scale alignment strategy effectively exploits both global scale cues and relative structural details, yielding the most accurate reconstruction and consistently more reliable results on 3D VL tasks.
Two-Stage Training. Fig. 6 compares different training strategies to assess their impact on convergence and performance. Training all modules jointly without staging (Single-stage) converges slowly and stabilizes at a relatively high loss value, due to gradient interference across modules. Using a single teacher under the two-stage strategy (Two-Stage (One-Teacher)) offers limited supervisory signals, resulting in less effective optimization. Removing the relational loss during two-stage strategy (Two-Stage (w/o RelaLoss)) further slows convergence and reduces accuracy, highlighting the importance of enforcing cross-task consistency. In contrast, using the full two-stage strategy with dual-teacher supervision and relational consistency (Two-Stage (Ours)) achieves the fastest and achieves the lowest final loss.